# Bayesian Optimization for Non-Convex Two-Stage Stochastic Optimization Problems

### v1

@@ 3.4 An alternative, alternating policy
The optimization of the joint knowledge gradient policy requires several layers of approximation. Previous works on robust (Pearce & Branke, 2017; Toscano-Palmerin & Frazier, 2022) and multi-task (Pearce & Branke, 2018) knowledge gradient which have used a discrete approximation for the inner maximization have avoided the need for the quasi-Monte Carlo (qMC) approximation of the expectation over the next observation. The tricks used cannot be extended to our formulation of the joint KG acquisition function. However, we can implement it in an alternating strategy switching between a pair of acquisition functions that aim to improve the fixed design, ${\bm{x}}$, and adjustable variables, ${\bm{y}}$, separately. Both acquisition functions use a GP model of the objective as a function of the whole space ${\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}$, and rely on the knowledge gradient principle. Algorithm 3 outlines the algorithm, and the two acquisition functions are explained in detail in the following two sections.
@@
3:     Fit hyperparameters of GP prior, $f\sim\mathcal{GP}(\mu,k)$, as described in Section E.3
@@
14:  Optimize ${\bm{x}}^{*n_{\mathrm{tot}}}\in\operatorname*{arg\,max}_{\bm{x}}\frac{1}{N_{u,\text{rec}}}\sum_{{\bm{u}}\in{\mathbb{U}}_{\mathrm{MC},\text{rec}}}\max_{{\bm{y}}}\mu^{n_{\mathrm{tot}}}({\bm{x}},{\bm{y}},{\bm{u}})$
15:  Let ${\bm{g}}^{*n_{\mathrm{tot}}}:{\mathbb{U}}\to{\mathbb{Y}}$ given by ${\bm{g}}^{*n_{\mathrm{tot}}}({\bm{u}})\in\operatorname*{arg\,max}_{{\bm{y}}}\mu^{n_{\mathrm{tot}}}({\bm{x}}^{*n_{\mathrm{tot}}},{\bm{y}},{\bm{u}})$
15:  ${\bm{x}}^{*n_{\mathrm{tot}}},{\bm{g}}^{*n_{\mathrm{tot}}}$
@@ 3.4.1 Improving the fixed design
The acquisition function designed to improve the fixed design asks ‘if we must choose the control policy ${\bm{g}}$ now but can make one more observation before choosing the fixed design ${\bm{x}}$, how much will the next sample improve the maximum of the posterior mean, in expectation?’. While this is a less natural question than the one that motivates the joint strategy, it effectively reduces our problem to a standard stochastic optimization problem, for which we can apply similar techniques to existing robust knowledge gradient formulations using discretizations (Pearce & Branke, 2017; Toscano-Palmerin & Frazier, 2022). Note that when we recommend the final solution, we still use Equation 10, making use of all information available.
@@
$\alpha_{\text{aKG-fix}}^{n}(\tilde{{\bm{x}}})$ $=\mathbb{E}_{{\textnormal{v}}^{n+1}}\biggl{[}\;\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\bar{\mu}^{n+1}({\bm{x}}^{\prime},{\bm{g}}^{*n})\,\bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\biggr{]}$ $-\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\bar{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{*n})$ (14a)
$=\mathbb{E}_{{\textnormal{v}}^{n+1}}\Biggl{[}\;\underbrace{\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\,\mathbb{E}_{{\mathbf{u}}^{\prime}}\!\biggl{[}\mu^{n+1}({\bm{x}}^{\prime},{\bm{g}}^{*n}({\mathbf{u}}^{\prime}),{\mathbf{u}}^{\prime})\biggr{]}}_{\begin{subarray}{c}\text{Maximum of posterior mean}\\ \text{after next observation, }{\textnormal{v}}^{n+1},\\ \text{if we can only change }{\bm{x}}^{\prime}\end{subarray}}\,\Bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\Biggr{]}$ $-\underbrace{\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\,\mathbb{E}_{{\mathbf{u}}^{\prime}}\!\biggl{[}\mu^{n}({\bm{x}}^{\prime},{\bm{g}}^{*n}({\mathbf{u}}^{\prime}),{\mathbf{u}}^{\prime})\biggr{]}}_{\begin{subarray}{c}\text{Current maximum of posterior mean}\\ \text{(constant wrt }\tilde{{\bm{x}}}\text{)}\end{subarray}}.$ (14b)
@@
$\hat{{\bm{x}}}^{*n}$ $\in\operatorname*{arg\,max}_{{\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}}\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\max_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu^{n}({\bm{x}}^{\prime},{\bm{y}}^{\prime},{\bm{u}}^{\prime}),$ (15a)
$\forall\,{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}\quad\hat{{\bm{g}}}^{*n}({\bm{u}}^{\prime})$ $\in\operatorname*{arg\,max}_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu^{n}(\hat{{\bm{x}}}^{*n},{\bm{y}}^{\prime},{\bm{u}}^{\prime}),$ (15b)
$\begin{split}\hat{\alpha}_{\text{aKG-fix}}^{n}(\tilde{{\bm{x}}})&=\mathbb{E}_{{\textnormal{v}}^{n+1}}\Biggl{[}\;\max_{{\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}}\underbrace{\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\mu^{n+1}({\bm{x}}^{\prime},\hat{{\bm{g}}}^{*n}({\bm{u}}^{\prime}),{\bm{u}}^{\prime})}_{\text{Varies linearly with }{\textnormal{v}}^{n+1}}\,\Bigg{|}\,\tilde{{\bm{x}}}^{n+1}={\bm{x}}\,\Biggr{]}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\mu^{n}({\bm{x}}^{\prime},\hat{{\bm{g}}}^{*n}({\bm{u}}^{\prime}),{\bm{u}}^{\prime}).\end{split}$ (15c)
@@ 3.4.2 Improving the adjustable variables
$\alpha_{\text{aKG-adj}}^{n}(\tilde{{\bm{x}}})$ $=\mathbb{E}_{{\textnormal{v}}^{n+1}}\biggl{[}\max_{{\bm{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}}\bar{\mu}^{n+1}({\bm{x}}^{*n},{\bm{g}}^{\prime})\,\bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\biggr{]}$ $-\max_{{\bm{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}}\bar{\mu}^{n}({\bm{x}}^{*n},{\bm{g}}^{\prime})\bigr{]}$ (16a)
$=\mathbb{E}_{{\textnormal{v}}^{n+1}}\Biggl{[}\;\underbrace{\mathbb{E}_{{\mathbf{u}}^{\prime}}\!\biggl{[}\max_{{\bm{y}}^{\prime}\in{\mathbb{Y}}}\mu^{n+1}({\bm{x}}^{*n},{\bm{y}}^{\prime},{\mathbf{u}}^{\prime})\biggr{]}}_{\begin{subarray}{c}\text{Maximum of posterior mean}\\ \text{after next observation, }{\textnormal{v}}^{n+1},\\ \text{if we can only change }{\bm{y}}^{\prime}\end{subarray}}\,\Bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\Biggr{]}$ $-\underbrace{\mathbb{E}_{{\mathbf{u}}^{\prime}}\!\biggl{[}\max_{{\bm{y}}^{\prime}\in{\mathbb{Y}}}\mu^{n}({\bm{x}}^{*n},{\bm{y}}^{\prime},{\mathbf{u}}^{\prime})\biggr{]}}_{\begin{subarray}{c}\text{Current maximum of posterior}\\ \text{mean (constant wrt }{\bm{x}},{\bm{y}}\text{ and }{\bm{u}}\text{)}\end{subarray}}.$ (16b)
@@
$\hat{\alpha}_{\text{aKG-adj}}^{n}(\tilde{{\bm{x}}})=\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\ \mathbb{E}_{{\textnormal{v}}^{n+1}}\Biggl{[}\max_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\!\!\underbrace{\mu^{n+1}(\hat{{\bm{x}}}^{*n},{\bm{y}}^{\prime},{\bm{u}}^{\prime})}_{\text{Varies linearly with }{\textnormal{v}}^{n+1}}\,\Bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\,\Biggr{]}\\ -\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\max_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu^{n}(\hat{{\bm{x}}}^{*n},{\bm{y}}^{\prime},{\bm{u}}^{\prime})$ (17)
where $\hat{{\bm{x}}}^{*n}$ is given by Equation 15a. Again, the expectation over ${\textnormal{v}}^{n+1}$ can be computed exactly, the result is differentiable with respect to $\tilde{{\bm{x}}}$, and the acquisition function can be optimized using a gradient-based optimizer such as multi-start L-BFGS-B.
@@ 3.5 The two-step incumbent method
In the absence of co-design, the fixed design ${\bm{x}}$ and controller ${\bm{g}}$ must be designed separately. That is, engineers must alternate between the following two steps:
@@
This is inefficient because it cannot be known up-front for how long to optimize each part of the design. For this reason, we consider a benchmark algorithm that only executes each step once and consider that future iterations of the steps would happen beyond the evaluation budget considered.
As in the previous sections, we employ the knowledge gradient acquisition function. In the first step, an ${\bm{x}}_{\text{step-1}}\in{\mathbb{X}}$ is chosen, the restriction of $h$ to $\{{\bm{x}}_{\text{step-1}}\}\times{\mathbb{Y}}\times{\mathbb{U}}$ is modeled using a Gaussian process, $f_{\text{step-1}}:{\mathbb{Y}}\times{\mathbb{U}}\to\mathbb{R}$, and the knowledge gradient is maximized to select the ${\bm{y}}$ and ${\bm{u}}$ for the next observation. At the end of the first step, an optimal policy ${\bm{g}}^{*n_{\mathrm{tot},1}}$ is extracted. In the second step, the restriction of $h$ to $\{({\bm{x}},{\bm{y}},{\bm{u}})\in{\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}:{\bm{y}}={\bm{g}}^{*n}({\bm{u}})\}$ is modeled with a Gaussian process, $f_{\text{step-2}}:{\mathbb{X}}\times{\mathbb{U}}\to\mathbb{R}$, and the knowledge gradient is maximized to select the ${\bm{x}}$ and ${\bm{u}}$ for the next observation. At the end of the second step, an optimal fixed design ${\bm{x}}^{*n_{\mathrm{tot},2}}$ is extracted.
@@
$\alpha_{\text{2sKG-1}}^{n}({\bm{y}},{\bm{u}})$ $=\mathbb{E}_{{\textnormal{v}}_{\text{step-1}}^{n+1}}\Biggl{[}\;\mathbb{E}_{{\mathbf{u}}^{\prime}}\biggl{[}\max_{{\bm{y}}^{\prime}\in{\mathbb{Y}}}\mu_{\text{step-1}}^{n+1}({\bm{y}}^{\prime},{\mathbf{u}}^{\prime})\biggr{]}\,\Bigg{|}\,({\bm{y}}_{\text{step-1}}^{n+1},{\bm{u}}_{\text{step-1}}^{n+1})=({\bm{y}},{\bm{u}})\Biggr{]}-\mathbb{E}_{{\mathbf{u}}^{\prime}}\biggl{[}\max_{{\bm{y}}^{\prime}\in{\mathbb{Y}}}\mu_{\text{step-1}}^{n}({\bm{y}}^{\prime},{\mathbf{u}}^{\prime})\biggr{]},$ (18a)
$\alpha_{\text{2sKG-2}}^{n}({\bm{x}},{\bm{u}})$ $=\mathbb{E}_{{\textnormal{v}}_{\text{step-2}}^{n+1}}\Biggl{[}\;\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\mathbb{E}_{{\mathbf{u}}^{\prime}}\bigl{[}\mu_{\text{step-2}}^{n+1}({\bm{x}}^{\prime},{\mathbf{u}}^{\prime})\bigr{]}\,\Bigg{|}\,({\bm{x}}_{\text{step-2}}^{n+1},{\bm{u}}_{\text{step-2}}^{n+1})=({\bm{x}},{\bm{u}})\Biggr{]}-\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\mathbb{E}_{{\mathbf{u}}^{\prime}}\bigl{[}\mu_{\text{step-2}}^{n}({\bm{x}}^{\prime},{\mathbf{u}}^{\prime})\bigr{]}.$ (18b)
@@
$\begin{split}\hat{\alpha}_{\text{2sKG-1}}^{n}({\bm{y}},{\bm{u}})&=\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\mathbb{E}_{{\textnormal{v}}_{\text{step-1}}^{n+1}}\Biggl{[}\;\max_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu_{\text{step-1}}^{n+1}({\bm{y}}^{\prime},{\bm{u}}^{\prime})\,\Bigg{|}\,({\bm{y}}_{\text{step-1}}^{n+1},{\bm{u}}_{\text{step-1}}^{n+1})=({\bm{y}},{\bm{u}})\Biggr{]}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\max_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu_{\text{step-1}}^{n}({\bm{y}}^{\prime},{\bm{u}}^{\prime}),\end{split}$ (19a)
$\begin{split}\hat{\alpha}_{\text{2sKG-2}}^{n}({\bm{x}},{\bm{u}})&=\mathbb{E}_{{\textnormal{v}}_{\text{step-2}}^{n+1}}\Biggl{[}\;\max_{{\bm{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}}\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\mu_{\text{step-2}}^{n+1}({\bm{x}}^{\prime},{\bm{u}}^{\prime})\,\Bigg{|}\,({\bm{x}}_{\text{step-2}}^{n+1},{\bm{u}}_{\text{step-2}}^{n+1})=({\bm{x}},{\bm{u}})\Biggr{]}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\max_{{\bm{x}}^{\prime},\in{{\mathbb{X}}_{\mathrm{disc}}}}\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\mu_{\text{step-2}}^{n}({\bm{x}}^{\prime},{\bm{u}}^{\prime}).\end{split}$ (19b)
@@
2:  Choose ${\bm{x}}_{\text{step-1}}\in{\mathbb{X}}$
@@
4:  for $n=n_{1}$ to $n_{\mathrm{tot},1}$ do
5:     Fit MAP hyperparameters of GP prior, $f_{\text{step-1}}\sim\mathcal{GP}$, on functions ${\mathbb{Y}}\times{\mathbb{U}}\to\mathbb{R}$ as described in Section E.3
@@
10:  Fit MAP hyperparameters of GP prior, $f_{\text{step-1}}\sim\mathcal{GP}$, on functions ${\mathbb{Y}}\times{\mathbb{U}}\to\mathbb{R}$ as described in Section E.3
11:  Let ${\bm{g}}^{*n_{\mathrm{tot},1}}:{\mathbb{U}}\to{\mathbb{Y}}$ given by ${\bm{g}}^{*n_{\mathrm{tot},1}}({\bm{u}})\in\operatorname*{arg\,max}_{{\bm{y}}}\mu_{\text{step-1}}^{n_{\mathrm{tot},1}}({\bm{y}},{\bm{u}})$
@@
13:  Evaluate $h$ at $n_{2}$ points, chosen using scrambled Sobol’ sequence on ${\mathbb{X}}\times{\mathbb{U}}$ augmented by ${\bm{y}}$-values given by ${\bm{g}}^{*n_{\mathrm{tot},1}}(\cdot)$
14:  for $n=n_{2}$ to $n_{\mathrm{tot},2}$ do
15:     Fit MAP hyperparameters of GP prior, $f_{\text{step-2}}\sim\mathcal{GP}$, on functions ${\mathbb{X}}\times{\mathbb{U}}\to\mathbb{R}$ as described in Section E.3
@@
18:     Evaluate $h\bigl{(}{\bm{x}}_{\text{step-2}}^{n+1},\,{\bm{g}}^{*n_{\mathrm{tot},1}}({\bm{u}}_{\text{step-2}}^{n+1}),\,{\bm{u}}_{\text{step-2}}^{n+1}\bigr{)}$ (expensive)

### v2

@@ 3.3 An alternative, alternating policy
The optimization of the joint knowledge gradient policy requires several layers of approximation. Previous works on robust (Pearce & Branke, 2017; Toscano-Palmerin & Frazier, 2022) and multi-task (Pearce & Branke, 2018) knowledge gradient, which used a discrete approximation for the inner maximization, have avoided the need for the quasi-Monte Carlo (qMC) approximation of the expectation over the next observation. The tricks used cannot be extended to our formulation of the joint KG acquisition function. However, we can implement it in an alternating strategy switching between a pair of acquisition functions that aim to improve the fixed design, ${\bm{x}}$, and adjustable variables, ${\bm{y}}$, separately. Both acquisition functions use a GP model of the objective as a function of the whole space ${\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}$, and rely on the knowledge gradient principle. Algorithm 3 outlines the algorithm, and the two acquisition functions are explained in detail in the following two sections.
@@
3:     Fit hyperparameters of GP prior, $f\sim\mathcal{GP}(\mu,k)$, using MAP estimates
@@
14:  Optimize $\hat{{\bm{x}}}^{*n_{\mathrm{tot}}}\in\operatorname*{arg\,max}_{\bm{x}}\frac{1}{N_{u,\text{rec}}}\sum_{{\bm{u}}\in{\mathbb{U}}_{\mathrm{MC},\text{rec}}}\max_{{\bm{y}}}\mu^{n_{\mathrm{tot}}}({\bm{x}},{\bm{y}},{\bm{u}})$
15:  Let $\hat{{\bm{g}}}^{*n_{\mathrm{tot}}}:{\mathbb{U}}\to{\mathbb{Y}}$ given by $\hat{{\bm{g}}}^{*n_{\mathrm{tot}}}({\bm{u}})\in\operatorname*{arg\,max}_{{\bm{y}}}\mu^{n_{\mathrm{tot}}}(\hat{{\bm{x}}}^{*n_{\mathrm{tot}}},{\bm{y}},{\bm{u}})$
15:  $\hat{{\bm{x}}}^{*n_{\mathrm{tot}}},\hat{{\bm{g}}}^{*n_{\mathrm{tot}}}$
@@ 3.3.1 Improving the fixed design
The acquisition function designed to improve the fixed design asks ‘if we must choose the control policy ${\bm{g}}$ now but can make one more observation before choosing the fixed design ${\bm{x}}$, how much will the next sample improve the maximum of the posterior mean, in expectation?’. While this is a less natural question than the one that motivates the joint strategy, it effectively reduces our problem to a standard stochastic optimization problem, for which we can apply similar techniques to existing robust knowledge gradient formulations using discretizations (Pearce & Branke, 2017; Toscano-Palmerin & Frazier, 2022). Note that when we recommend the final solution, we still use Equations 10 and 13, making use of all information available.
@@
$\alpha_{\text{aKG-fix}}^{n}(\tilde{{\bm{x}}})$ $=\mathbb{E}_{{\textnormal{v}}^{n+1}}\biggl{[}\;\max_{{\mathbf{x}}^{\prime}\in{\mathbb{X}}}\bar{\mu}^{n+1}({\mathbf{x}}^{\prime},{\bm{g}}^{*n})\,\bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\biggr{]}$ $-\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\bar{\mu}^{n}({\bm{x}}^{\prime},{\bm{g}}^{*n})$ (14a)
$=\mathbb{E}_{{\textnormal{v}}^{n+1}}\Biggl{[}\;\underbrace{\max_{{\mathbf{x}}^{\prime}\in{\mathbb{X}}}\,\mathbb{E}_{{\mathbf{u}}^{\prime}}\!\biggl{[}\mu^{n+1}({\mathbf{x}}^{\prime},{\bm{g}}^{*n}({\mathbf{u}}^{\prime}),{\mathbf{u}}^{\prime})\biggr{]}}_{\begin{subarray}{c}\text{Maximum of posterior mean}\\ \text{after next observation, }{\textnormal{v}}^{n+1},\\ \text{if we can only change }{\mathbf{x}}^{\prime}={\mathbf{x}}^{*(n+1)}\end{subarray}}\,\Bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\Biggr{]}$ $-\underbrace{\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\,\mathbb{E}_{{\mathbf{u}}^{\prime}}\!\biggl{[}\mu^{n}({\bm{x}}^{\prime},{\bm{g}}^{*n}({\mathbf{u}}^{\prime}),{\mathbf{u}}^{\prime})\biggr{]}}_{\begin{subarray}{c}\text{Current maximum of posterior mean}\\ \text{(constant wrt }\tilde{{\bm{x}}}\text{)}\end{subarray}}.$ (14b)
@@
[…]
@@ 3.3.2 Improving the adjustable variables
$\alpha_{\text{aKG-adj}}^{n}(\tilde{{\bm{x}}})$ $=\mathbb{E}_{{\textnormal{v}}^{n+1}}\biggl{[}\max_{{\mathbf{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}}\bar{\mu}^{n+1}({\bm{x}}^{*n},{\mathbf{g}}^{\prime})\,\bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\biggr{]}$ $-\max_{{\bm{g}}^{\prime}:{\mathbb{U}}\to{\mathbb{Y}}}\bar{\mu}^{n}({\bm{x}}^{*n},{\bm{g}}^{\prime})$ (16a)
$=\mathbb{E}_{{\textnormal{v}}^{n+1}}\Biggl{[}\;\underbrace{\mathbb{E}_{{\mathbf{u}}^{\prime}}\!\biggl{[}\max_{{\mathbf{y}}^{\prime}\in{\mathbb{Y}}}\mu^{n+1}({\bm{x}}^{*n},{\mathbf{y}}^{\prime},{\mathbf{u}}^{\prime})\biggr{]}}_{\begin{subarray}{c}\text{Maximum of posterior mean}\\ \text{after next observation, }{\textnormal{v}}^{n+1},\\ \text{if we can only change }{\mathbf{y}}^{\prime}={\mathbf{g}}^{*(n+1)}({\mathbf{u}}^{\prime})\end{subarray}}\,\Bigg{|}\,\tilde{{\bm{x}}}^{n+1}=\tilde{{\bm{x}}}\Biggr{]}$ $-\underbrace{\mathbb{E}_{{\mathbf{u}}^{\prime}}\!\biggl{[}\max_{{\mathbf{y}}^{\prime}\in{\mathbb{Y}}}\mu^{n}({\bm{x}}^{*n},{\mathbf{y}}^{\prime},{\mathbf{u}}^{\prime})\biggr{]}}_{\begin{subarray}{c}\text{Current maximum of posterior}\\ \text{mean (constant wrt }{\bm{x}},{\bm{y}}\text{ and }{\bm{u}}\text{)}\end{subarray}}.$ (16b)
@@
[…]
@@ 3.4 The two-step incumbent method
The standard approach of a particular industry partner is to design the fixed design ${\bm{x}}$ and controller ${\bm{g}}$ separately. That is, engineers must alternate between the following two steps:
@@
This is inefficient because there is no immediate convergence criterion to determine when to switch steps, and because the GP model must ‘forget’ past data each time it does switch steps. For this reason, we consider a benchmark algorithm that only executes each step once and consider that future iterations of the steps would happen beyond the evaluation budget.
As in the previous sections, we employ the knowledge gradient acquisition function. In the first step, an ${\bm{x}}_{\text{step-1}}\in{\mathbb{X}}$ is chosen arbitrarily, the restriction of $h$ to $\{{\bm{x}}_{\text{step-1}}\}\times{\mathbb{Y}}\times{\mathbb{U}}$ is modeled using a Gaussian process $f_{\text{step-1}}:{\mathbb{Y}}\times{\mathbb{U}}\to\mathbb{R}$ and the knowledge gradient (named REVI in (Pearce & Branke, 2018)) is maximized to select the ${\bm{y}}$ and ${\bm{u}}$ for the next observation. At the end of the first step, an optimal policy ${\bm{g}}^{*n_{\mathrm{tot},1}}$ is extracted. In the second step, the restriction of $h$ to $\{({\bm{x}},{\bm{y}},{\bm{u}})\in{\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}:{\bm{y}}={\bm{g}}^{*n_{\mathrm{tot},1}}({\bm{u}})\}$ is modeled with a Gaussian process $f_{\text{step-2}}:{\mathbb{X}}\times{\mathbb{U}}\to\mathbb{R}$ and the knowledge gradient is maximized to select the ${\bm{x}}$ and ${\bm{u}}$ for the next observation. At the end of the second step, an optimal fixed design ${\bm{x}}^{*n_{\mathrm{tot},2}}$ is extracted.
@@
$\alpha_{\text{2sKG-1}}^{n}({\bm{y}},{\bm{u}})$ $=\mathbb{E}_{{\textnormal{v}}_{\text{step-1}}^{n+1}}\Biggl{[}\;\mathbb{E}_{{\mathbf{u}}^{\prime}}\biggl{[}\max_{{\mathbf{y}}^{\prime}\in{\mathbb{Y}}}\mu_{\text{step-1}}^{n+1}({\mathbf{y}}^{\prime},{\mathbf{u}}^{\prime})\biggr{]}\,\Bigg{|}\,({\bm{y}}_{\text{step-1}}^{n+1},{\bm{u}}_{\text{step-1}}^{n+1})=({\bm{y}},{\bm{u}})\Biggr{]}-\mathbb{E}_{{\mathbf{u}}^{\prime}}\biggl{[}\max_{{\mathbf{y}}^{\prime}\in{\mathbb{Y}}}\mu_{\text{step-1}}^{n}({\mathbf{y}}^{\prime},{\mathbf{u}}^{\prime})\biggr{]},$ (18a)
$\alpha_{\text{2sKG-2}}^{n}({\bm{x}},{\bm{u}})$ $=\mathbb{E}_{{\textnormal{v}}_{\text{step-2}}^{n+1}}\Biggl{[}\;\max_{{\mathbf{x}}^{\prime}\in{\mathbb{X}}}\mathbb{E}_{{\mathbf{u}}^{\prime}}\bigl{[}\mu_{\text{step-2}}^{n+1}({\mathbf{x}}^{\prime},{\mathbf{u}}^{\prime})\bigr{]}\,\Bigg{|}\,({\bm{x}}_{\text{step-2}}^{n+1},{\bm{u}}_{\text{step-2}}^{n+1})=({\bm{x}},{\bm{u}})\Biggr{]}-\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\mathbb{E}_{{\mathbf{u}}^{\prime}}\bigl{[}\mu_{\text{step-2}}^{n}({\bm{x}}^{\prime},{\mathbf{u}}^{\prime})\bigr{]}.$ (18b)
@@
$\begin{split}\hat{\alpha}_{\text{2sKG-1}}^{n}({\bm{y}},{\bm{u}})&=\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\mathbb{E}_{{\textnormal{v}}_{\text{step-1}}^{n+1}}\Biggl{[}\;\max_{{\mathbf{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu_{\text{step-1}}^{n+1}({\mathbf{y}}^{\prime},{\bm{u}}^{\prime})\,\Bigg{|}\,({\bm{y}}_{\text{step-1}}^{n+1},{\bm{u}}_{\text{step-1}}^{n+1})=({\bm{y}},{\bm{u}})\Biggr{]}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\max_{{\bm{y}}^{\prime}\in{{\mathbb{Y}}_{\mathrm{disc}}}}\mu_{\text{step-1}}^{n}({\bm{y}}^{\prime},{\bm{u}}^{\prime}),\end{split}$ (19a)
$\begin{split}\hat{\alpha}_{\text{2sKG-2}}^{n}({\bm{x}},{\bm{u}})&=\mathbb{E}_{{\textnormal{v}}_{\text{step-2}}^{n+1}}\Biggl{[}\;\max_{{\mathbf{x}}^{\prime}\in{{\mathbb{X}}_{\mathrm{disc}}}}\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\mu_{\text{step-2}}^{n+1}({\mathbf{x}}^{\prime},{\bm{u}}^{\prime})\,\Bigg{|}\,({\bm{x}}_{\text{step-2}}^{n+1},{\bm{u}}_{\text{step-2}}^{n+1})=({\bm{x}},{\bm{u}})\Biggr{]}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\max_{{\bm{x}}^{\prime},\in{{\mathbb{X}}_{\mathrm{disc}}}}\frac{1}{N_{u}}\sum_{{\bm{u}}^{\prime}\in{{\mathbb{U}}_{\mathrm{MC}}}}\mu_{\text{step-2}}^{n}({\bm{x}}^{\prime},{\bm{u}}^{\prime}).\end{split}$ (19b)
@@
2:  Choose ${\bm{x}}_{\text{step-1}}\in{\mathbb{X}}$ arbitrarily
@@
4:  for $n=n_{1}$ to $n_{\mathrm{tot},1}-1$ do
5:     Fit MAP hyperparameters of GP prior, $f_{\text{step-1}}\sim\mathcal{GP}$, on functions ${\mathbb{Y}}\times{\mathbb{U}}\to\mathbb{R}$
@@
10:  Fit MAP hyperparameters of GP prior, $f_{\text{step-1}}\sim\mathcal{GP}$, on functions ${\mathbb{Y}}\times{\mathbb{U}}\to\mathbb{R}$
11:  Let $\hat{{\bm{g}}}^{*n_{\mathrm{tot},1}}:{\mathbb{U}}\to{\mathbb{Y}}$ given by $\hat{{\bm{g}}}^{*n_{\mathrm{tot},1}}({\bm{u}})\in\operatorname*{arg\,max}_{{\bm{y}}}\mu_{\text{step-1}}^{n_{\mathrm{tot},1}}({\bm{y}},{\bm{u}})$
@@
13:  Evaluate $h$ at $n_{2}$ points, chosen using scrambled Sobol’ sequence on ${\mathbb{X}}\times{\mathbb{U}}$ augmented by ${\bm{y}}$-values given by $\hat{{\bm{g}}}^{*n_{\mathrm{tot},1}}(\cdot)$
14:  for $n=n_{2}$ to $n_{\mathrm{tot},2}-1$ do
15:     Fit MAP hyperparameters of GP prior, $f_{\text{step-2}}\sim\mathcal{GP}$, on functions ${\mathbb{X}}\times{\mathbb{U}}\to\mathbb{R}$
@@
18:     Evaluate $h\bigl{(}{\bm{x}}_{\text{step-2}}^{n+1},\,\hat{{\bm{g}}}^{*n_{\mathrm{tot},1}}({\bm{u}}_{\text{step-2}}^{n+1}),\,{\bm{u}}_{\text{step-2}}^{n+1}\bigr{)}$ (expensive)